The demand equation for a particular kind of shirt is 2px+65p-4950=0, where x hundreds of shirts are demanded per week when dollars is the price of a shirt. Suppose that the shirt is selling this week at $30, and the price increases at the rate of $0.20 per week. Find the rate of change in demand. 


1
Expert's answer
2021-07-26T16:09:19-0400
"2px+65p-4950=0"

Differentiate both sides with respect to "t"


"2p'x+2px'+65p'=0"

Solve for "x'"


"x'=-\dfrac{p'(2x+65)}{2p}"

Given "p=30, p'=0.20"


"2(30)x+65(30)-4950=0"

"x=50"

Substitute


"x'=-\dfrac{0.2(2(50)+65)}{2(30)}=-1.1"

 The rate of change in demand is "-110" shirts per week.